You will use 1.00microgram of each plasmid in a given transformation reaction. The plasmids are at the concentrations listed below. For each of the six different plasmids, how many microliters equals 1.00 micrograms?

#115: .910mg/ml

#138: .558mg/ml

#168: .918mg/ml

#140: .668mg/ml

#81: .430mg/ml

#51: 1.385mg/ml
Answer #1

#115:    910ug ------------ 1ul

           1 ug -------------- 1 X 1 / 0.910 = 1.09 ul

#138: .558ug ------------ 1 ul

             1 ug ------------ 1 X 1 / 0.558 = 1.79 ul

#168: .918mg/ml

              .918ug ------------ 1ul

             1 ug ------------ 1 X 1 / 0.918 = 1.08 ul

#140: .668mg/ml

               .668ug ------------ 1ul

               1 ug ------------ 1 X 1 / 0.668 = 1.49 ul

#81: .430mg/ml

                .430ug ------------ 1ul

                1 ug ------------ 1 X 1 / 0.430 = 2.32 ul

#51: 1.385mg/ml

                  1.385ug ------------ 1ul

                 1 ug ------------ 1 X 1 / 1.385 = 0.72 ul